Jane works due north of home.her husband ben works due east.they leave for work at the same time.by the time jane is 5 miles from home,the distance between them is one mile more than ben's distance. how far is ben from home?
Answer by ankor@dixie-net.com(22740)

Jane works due north of home.
her husband ben works due east.they leave for work at the same time.
by the time jane is 5 miles from home,the distance between them is one mile more than ben's distance.
how far is ben from home?
:
let x = Ben's distance from home
then
(x+1) = the distance between them when J is 5 miles from home
:
This is a pythag problems a^2 + b^2 = c^2, where
a = 5, b = x, c =(x+1)
:
5^2 + x^2 = (x+1)^2
25 + x^2 = x^2 + 2x + 1
subtract x^2 from both sides
25 = 2x + 1
25 - 1 = 2x
24 = 2x
x = 12 mi is Ben's distance from home
:
:
Check this out:
25 + 144 = 169 which is 13^2